# Caiqi Zhang1, Fangyu Liu1, Marco Basaldella2, Nigel Collier1

### v1

@@ 4.2 Baselines for UQ
4.3 Uncertainty Quantification Results
FS PFS US PUS RR
GPT-4 80.76 72.37 20.75 28.98 86.62
GPT-3.5 68.25 67.31 25.71 25.71 100
Yi-34B-Chat 55.71 55.71 41.25 41.25 100
Tulu-2-70B 47.19 47.19 55.83 55.83 100
Gemini Pro 43.20 42.73 61.74 62.17 98.90
Vicuna-33B 42.47 42.47 55.31 55.31 100
Table 2: Results on FactScore dataset; FS and PFS stand for average factuality score and penalized factuality score; US and PUS stand for average uncertainty score and penalized uncertainty score computed by MutualNLI; RR means the respond rate. All numbers are in percentages.
(a) GPT-4
(b) GPT-3.5
(c) Gemini Pro
(d) Yi-34B-Chat
(e) Tulu-2-70B
(f) Vicuna-33B
Effectiveness of Luq
Table 1 and Figure 2 illustrate the correlation between factuality scores and uncertainty scores. The results highlight Luq’s effectiveness as an indicator of model factuality in long text generation tasks. Luq demonstrates a strong negative correlation for GPT-3.5, Gemini Pro, Yi-34B-Chat, Vicuna-33B, and Tulu-2-70B, with the strongest Pearson correlation being -0.8509.
For the baseline methods, LexSimilarity emerges as a robust baseline offering lower computational demands. EigValLaplacian have competing performance with LexSimilarity in models like Gemini Pro and Vicuna-33B. The confidence-based SelfCheckNLI method demonstrates the best Spearman correlation in models such as Gemini Pro and Tulu-2-70B. Other baselines such as Eccentricity, NumSemSets and DegMat yield unsatisfactory results, occasionally exhibiting even positive correlations.
We also observe that Luq is better suited for models with relatively lower factuality and a lack of self-awareness regarding uncertainty. For models with high factuality capabilities, such as GPT-4, Luq only demonstrates a moderate correlation with factuality scores. As indicated in Table 2 and Figure 2(a), among all models, GPT-4 exhibits the highest overall factuality scores and the lowest average uncertainty scores. The data points are tightly clustered with only few instances of uncertainty. Moreover, GPT-4’s self-awareness of uncertainty may also impact external UQ efficiency, as it tends to abstain from answering questions more often compared to other models, highlighting its heightened internal uncertainty awareness. This observation does not influence the effectiveness of our method, as in real life models with lower factuality and unable to express uncertainty are in greater need of external uncertainty measurements.
Higher frequency leads to higher factuality and lower uncertainty.
In Figure 3, we compare the factuality and uncertainty scores across different entity frequencies. The original FActScore dataset provides the frequency of each entity in Wikipedia, categorizing them based on page views and co-occurrence within the training set (Min et al., 2023). Frequencies are classified into five categories, ranging from “very rare” to “very frequent.” Our observations suggest that questions associated with higher entity frequencies tend to yield more factual responses, alongside decreased model uncertainty. Notably, GPT-4 demonstrates consistent performance regarding uncertainty and factuality across varying frequencies, potentially attributable to its selective response strategy. Although it answers all the questions in the “very frequent,” “frequent,” and “medium” categories, it refuses to answer around 25% of “rare” questions and 30% of “very rare” questions.
Figure 3: Comparison of factuality and uncertainty scores across different frequencies.
GPT-3.5 Yi-34B-Chat Tulu-2-70B Vicuna-33B Gemini-Pro
Percentile FS US FS US FS US FS US FS US
0 68.25 25.71 55.71 41.25 47.19 55.83 42.47 55.31 43.20 61.74
2.5 69.80 24.07 56.91 40.24 48.28 53.86 43.59 54.41 44.28 60.05
5 70.82 23.40 57.99 39.28 49.35 53.13 44.54 53.76 45.20 59.23
7.5 71.52 22.66 58.86 38.12 50.27 52.57 45.54 52.97 46.34 58.19
10 72.30 22.18 60.21 36.79 51.45 51.87 46.14 52.34 47.25 57.73
12.5 74.13 21.61 61.69 35.04 52.13 51.31 46.53 51.58 48.41 56.29
15 75.04 21.17 62.91 34.18 53.29 50.63 47.52 50.96 49.51 55.40
Table 3: Selective question answering results. The percentile indicates the percentage of questions for which answers were abstained. All numbers are expressed as percentages.
Methods PFS PUS AD
Tulu-2-70B 47.19 55.83 42.08
Gemini Pro 42.73 62.17 29.51
Vicuna-33B 42.47 58.06 28.41
Luq-Ensemble 52.83 45.83 100
Yi-34B-Chat 55.71 41.25 66.12
Tulu-2-70B 47.19 55.83 21.31
Gemini Pro 42.73 62.17 12.57
Luq-Ensemble 58.75 37.60 100
GPT-3.5 67.31 25.71 92.35
Gemini Pro 42.73 62.17 1.64
Vicuna-33B 42.47 58.06 6.01
Luq-Ensemble 67.37 24.75 100
GPT-4 72.11 28.98 60.11
GPT-3.5 67.31 25.71 32.79
Yi-34B-Chat 55.71 41.25 7.10
Luq-Ensemble 76.61 17.27 100
Table 4: Results of different ensemble strategies. The Answer Distribution (AD) metric indicates the percentage of final answers generated by each model in the ensemble. All numbers are presented in percentages.
4.4 Luq-Ensemble
Given the variance in training corpus, different LLMs may possess varying levels of knowledge for a specific question, leading to discrepancies in uncertainty levels across models. Utilizing the Luq uncertainty score as a reliable indicator of factuality, we enhance overall performance through an ensemble approach. In this method, the model exhibiting the lowest Luq score for a given question is chosen as the final answer. Experimental results (Table 4) affirm the superiority of the Luq-Ensemble over its constituent counterparts.
Furthermore, we observe that ensembling models with similar factuality scores can notably enhance performance. For instance, in the combination of Tulu-2-70B, Gemini Pro, and Vicuna-33B, the PFS increases by 5% compared to the originally top-performing Tulu-2-70B, which scored 47.19%. Combining more comparable models also results in a more balanced answer distribution. In contrast, integrating a model with substantially superior performance, as seen in the combination of GPT-3.5, Gemini Pro, and Vicuna-33B, predominantly favors answers from GPT-3.5 (92.35%), leading to marginal improvement (0.06%) in the ensemble method.
To provide a more detailed analysis, Table 4 includes an ensemble of the top three models with the highest factuality scores: GPT-4, GPT-3.5, and Yi-34B-Chat. The results indicate that the majority of answers are generated by GPT-4 and GPT-3.5, accounting for 92.90%. However, the improvement in PFS primarily stems from questions that GPT-4 refuses to answer. They are given a factuality score of zero and an uncertainty score of one for penalty. This underscores a pivotal decision point for practitioners regarding model behavior in the face of uncertainty: whether the model should attempt to answer even though it is uncertain or opt to abstain.
4.5 Selective Question Answering
From Table 2, it is observed that while GPT-4 opts not to respond to some queries, other models generally attempt to answer all questions. The limited refusal by Gemini Pro primarily stems from considerations of sensitive content and regulatory constraints, rather than uncertainty [^1] Gemini Pro API returns the reasons of refusing to answer certain questions.. Therefore, we investigate the application of the Luq score to equip these models with the capability for selective question answering—that is, to enable them to decline responses when uncertain. Contrary to the traditional aim of responding correctly to every question, the objective in a selective question answering framework is to preserve accuracy while maximizing the number of questions answered (Kamath et al., 2020; Cole et al., 2023; Yang et al., 2023).
Table 3 presents the results of selective question answering. The models are permitted to refrain from answering questions with high uncertainty. The percentiles indicate the proportion of questions each model abstained from answering. The findings demonstrate that adopting a selective answering approach enhances the models’ factuality by allowing for more question rejections. By declining to answer a similar proportion of questions (approximately 15%) as GPT-4, the models typically achieve an improvement of over 5% in overall factuality scores. Notably, when implementing this selective answering strategy in practical applications, it is essential for practitioners to tailor the uncertainty thresholds to the specific models and tasks at hand.
4.6 Ablation Study
@@
Figure 4: The effect of different temperatures (upper) and the number of samples (lower) on the PCC with MutualNLI.
5 Related Work
UQ in Machine Learning
Prior to LLMs, UQ has been extensively explored within the field of machine learning (Gawlikowski et al., 2023). According to the source of uncertainty, it is typically categorized into two types: aleatoric and epistemic uncertainty(Hora, 1996; Der Kiureghian and Ditlevsen, 2009). Aleatoric uncertainty, also known as statistical uncertainty, pertains to the inherent randomness in experimental outcomes due to stochastic effects (Hüllermeier and Waegeman, 2021). In contrast, epistemic uncertainty stems from incomplete knowledge, potentially including uncertainties in a machine learning model’s parameters or the lack of certain training data (Hüllermeier and Waegeman, 2021; Huang et al., 2023). Our focus is primarily on epistemic uncertainty.
Gawlikowski et al. (2023) categorized uncertainty estimation into four types: (1) Single deterministic methods use a single pass in a deterministic network for predictions (Oberdiek et al., 2018; Lee and AlRegib, 2020); (2) Bayesian methods encompass various types of stochastic neural networks, where predictions are stochastic, each determined by differing model weights (Gal and Ghahramani, 2016; Welling and Teh, 2011; Lee et al., 2020); (3) Ensemble methods merge outputs from several deterministic networks, valuing model diversity (Wen et al., 2019; Lakshminarayanan et al., 2017); (4) Test-time augmentation methods generate multiple predictions through input data augmentation at test-time in a deterministic network (Lyzhov et al., 2020; Kim et al., 2020). Our method aligns with test-time augmentation but repeats identical questions at inference without data augmentation.
UQ in LLMs
In contrast to discriminative models, which readily provide probability scores for specific categories, uncertainty estimation in generative LLMs presents unique challenges: (1) There is an exponential increase in the output space as sentence length grows, rendering the evaluation of all possible predictions impractical (Geng et al., 2023; Wang et al., 2023). (2) The significance of semantic nuances and their inherent uncertainties, which diverges from the fixed category labels typical of discriminative models, complicates matters further (Kuhn et al., 2022). Generally, UQ methods for LLMs can be categorized based on the accessibility of the model’s internal states, distinguishing between black-box and white-box approaches. White-box LLMs often rely on logit-based evaluations, assessing sentence uncertainty through token-level probabilities or entropy (Murray and Chiang, 2018; Kuhn et al., 2022; Vazhentsev et al., 2023; Duan et al., 2023).
However, as access to LLMs increasingly relies on API calls, research has pivoted towards black-box methods. These can be further categorized into: (i) verbalized methods, which prompt LLMs to articulate their uncertainty in the output, using phrases like “I am sure" or “I do not know" (Mielke et al., 2022). Nonetheless, a practical mismatch between the expressed and actual uncertainty levels has been noted (Lin et al., 2022; Xiong et al., 2023). Xiong et al. (2023) highlight that LLMs often display excessive confidence when verbalizing their certainty. (ii) Consistency-based (sampling-based) estimation premises on the assumption that increased uncertainty in a model corresponds to greater diversity in its outputs, frequently resulting in hallucinatory outputs (Manakul et al., 2023; Lin et al., 2023). Our proposed method, MutualNLI, follows this consistency-based approach. There are also efforts on integrating verbalized methods with consistency-based approaches (Xiong et al., 2023; Rivera et al., 2024).
6 Conclusion
In this work, we first identify that existing UQ methods are ineffective on long text generation. We therefore introduce Luq, a novel UQ method tailored for long-form text generation in LLMs. It overcomes the limitation of previous methods by calculating sentence level consistency. We conduct extensive experiments over six popular LLMs, such as GPT-4 and Gemini Pro. Our findings demonstrate that Luq significantly improves the correlation with models’ factuality scores over previous methods. Luq serves as a reliable indicator of model’s factuality performance. Additionally, we present Luq-Ensemble, a model ensembling and selective question answering strategy, which showcases a promising avenue for enhancing the factual accuracy of LLM outputs. This research not only advances our understanding of UQ in the context of LLMs but also offers practical tools for improving the reliability and trustworthiness of AI-generated content.
Limitation
The limitations of this study include the following: (1) A primary challenge in studying uncertainty quantification for long text generation lies in the difficulty of evaluating the generated text. Unlike classification tasks and short-answer QA, there is no straightforward metric for assessing the quality of generated text. In this study, we employ the factuality score as the primary evaluation metric, thereby leaving other text aspects, such as coherence, cohesion, and creativity, under-explored. Future work could investigate uncertainty scores using more comprehensive evaluation metrics. (2) Fact-checking for long text generation is costly, especially when involving human annotation. While we utilized FActScore as an automated evaluation tool, it still requires numerous API calls to fact-check each piece of information. The development of more cost-effective evaluation metrics or datasets could allow for the expansion of the experimental scope. (3) We discovered that temperature plays a crucial role in measuring uncertainty. A common limitation among all consistency-based uncertainty quantification methods is their effectiveness only at high temperatures. In practice, if some scenarios necessitate a fixed low temperature, these consistency-based UQ methods may not perform effectively.
Ethics Statement
Our research strictly follows ethical guidelines, focusing on data privacy, bias mitigation, and societal impact. We use the publicly available FActScore dataset, which ensures a balanced representation of different nationalities. Our code usage complies with original licensing agreements and is strictly for academic purposes, reflecting our commitment to ethical research standards.
Appendix A Experiment Setup
For GPT-4 and GPT-3.5, we use the OpenAI API, with specific version gpt-4-turbo-0125-preview and gpt-3.5-turbo-0613. For Gemini Pro, we call the API for developers. For Yi-34B-Chat, Tulu-2-70B (tulu-2-dpo-70b), and Vicuna-33B (vicuna-33b-v1.3), we use them off-the-shelf and only for inference. We run our uncertainty measurement experiments on A100-SXM-80GB GPUs. For our experiments, we use the following prompt:
{mdframed}
Appendix B Number of Facts in a Response
Figure 5: Average Number of Atomic Facts (AF) in the Response
Figure 5 shows the average factual content provided by various AI models in the FactScore dataset. GPT-4 has the highest average of atomic facts at 52.24, indicating it provides the most detailed factual responses. Tulu-2-70B follows with an average of 52.17, nearly matching GPT-4 in factual detail. GPT-3.5 has an AF of 50.67, showing it also delivers a high level of factual detail in its responses. Yi-34B-Chat and Gemini Pro have lower averages, at 45.80 and 42.72 respectively. Vicuna-33B has the lowest AF at 36.20, indicating it offers the least amount of factual information in its responses. Generally, these models provide similar number of atomic facts in their responses.

### v2

@@ Appendix E Baselines
Appendix F Response Statistics
Average Response Length.
In our study, we define “long text” as the model’s output consisting of at least 100 words. For our experiments, the texts are even longer, averaging more than 200 words. The average word count is calculated and shown in the following table. In contrast, commonly used datasets for existing UQ methods have much shorter texts, with an average of 1.95 words for Trivia QA dataset and 3.37 words for Natural Questions dataset.
Model Avg. Word Count per Response
GPT-4-0125 264
GPT-3.5-turbo-0613 239
Gemini Pro 1.0 223
Yi-34B-Chat 307
Tulu-2-70B 267
Vicuna-33B 206
Table 9: The average response length for each LLM.
Number of Facts in a Response
Figure 4 shows the average atomic facts provided by various AI models for the FActScore dataset. GPT-4 has the highest average number of atomic facts at 52.24, indicating it provides the most detailed factual responses. Tulu-2-70B follows with an average of 52.17, nearly matching GPT-4 in factual details. GPT-3.5 has an AF of 50.67, showing it also delivers a high level of factual details in its responses. Yi-34B-Chat and Gemini 1.0 Pro have comparatively lower averages, at 45.80 and 42.72 respectively. Vicuna-33B has the lowest AF at 36.20, indicating it offers the least amount of factual information in its responses. Generally, these models provide similar number of atomic facts in their responses.
Figure 4: Average number of atomic Facts (AF) in a response for each model.
Appendix G Luq-Ensemble
In this section, we discuss more about the novelty, motivation, and effectiveness of our Luq-Ensemble method.
The Motivation of Luq-Ensemble. After getting multiple answers from different LLMs, the challenge is now to choose which one as the final output. Traditional aggregation methods like majority vote and weighted vote are ineffective for long text generation because finding the majority answer is difficult when all the responses are somewhat different, and methods like boosting or bagging require additional training. Our uncertainty measure thus serves as an effective indicator for model ensembling.
The Effectiveness of Luq-Ensemble. Meanwhile, it is not guaranteed that ensembling will lead to better performance. It is true that the uncertainty scores will by definition decrease (as we select the model with the least uncertainty as the final response), but the factuality score may not. The effectiveness of Luq-Ensemble largely relies on the reliability of the UQ method.
Table 10 compares the effectiveness of using Luq as the ensemble indicator with other methods. The results show that ensembling does not inherently improve performance. With a poor UQ method (such as Ecc), the ensemble factuality score can be lower than that of its highest component (47.2 vs 43.3). In contrast, using Luq as the indicator for ensembling yields the best overall performance.
PFS PUS
Individual Results
Tulu-2-70B 47.2 55.8
Gemini 1.0 Pro 42.7 62.2
Vicuna-33B 42.5 58.1
Different Ensemble Methods
Ecc-Ensemble 43.4 35.5
LexSim-Ensemble 47.6 39.8
SelfCheck-Ensemble 49.3 46.7
Luq-Ensemble 52.8 45.8
Table 10: Penalised factuality score (PFS) and Penalised uncertainty score (PUS) for individual models and ensembles with different UQ methods.
Appendix H Selective QA Strategy
When implementing a selective answering strategy in practical applications, it is essential for practitioners to tailor the uncertainty thresholds to the specific models and tasks at hand. In our experiment, as shown in Table 2, we find that GPT-4 tends to refuse to answer around 15% of the questions. To simulate a GPT-4-like answering strategy, for each model, we set different thresholds to ensure they refuse to answer between 0 and 15% of the questions. Our experiments indicate that different LLMs may have varying average absolute uncertainty values, making a universal uncertainty threshold unsuitable for all models. Additionally, the inherent nature of the tasks may influence practitioners’ decisions to make the model more conservative or more willing to attempt answering users’ questions.
We advise practitioners to implement selective QA strategies using the following practical steps:
  • Collect a representative set of questions/queries that closely mimic real-world usage scenarios.
  • Obtain responses from the LLMs for these questions and apply UQ methods (e.g., Luq) to get uncertainty scores.
  • Establish thresholds tailored to the specific task and the practitioners’ goals, selecting either cautious (lower threshold) or lenient (higher threshold) settings as needed.
  • Develop clear strategies for handling high uncertainty, such as refusing to answer, requesting clarification, or using alternative approaches.
Appendix I Ablation Study
@@
Figure 5: The effect of different temperatures (upper) and the number of samples (lower) on the PCC with Luq.
Appendix J Case Study
In this section, we present two case studies illustrating the performance of Luq. For simplicity, we only show three samples for each question using Yi-34B-Chat. In both cases, the baseline UQ methods indicate low uncertainty. However, the first case is with higher factuality, while the second case includes many non-factual information. Unlike baseline UQ methods that can only calculate similarity at the passage level, our method captures more nuanced differences.
In the first case study, Luq demonstrates lower uncertainty. As shown in Figure 6, we use green to represent sentences that can be cross-validated by other samples. It is evident that most sentences are consistent across different samples, thus leading to lower uncertainty.
In the second case study, we showcase an example with higher Luq uncertainty. In Figure 7, red is used to indicate contradictory information. It is apparent that the samples are not consistent with each other. The samples even state different nationalities of Jonathan Haagensen.
Figure 6: Example with low uncertainty by Luq. Sentences that can be cross-validated by other samples are shown in green.
Figure 7: Example with high uncertainty by Luq. Contradictory information is highlighted in red.